OpenStudy (anonymous):

The radius of a circle is 9cm. By what amount must the radius be increased to increase the area by 19π square centimeters? (HINT: Recall that the area of a circle is A = πr2 .)

OpenStudy (anonymous):

\[A=\pi r^2\] The radius is 9, so you have \[A=9^2\pi=81\pi\] You want area to increase by 19π, so you want to solve for x in this equation: \[81\pi+19\pi=81\pi+(9+x)^2\pi\\ (81\pi+19\pi=\text{Previous area $+$ new area})\]

OpenStudy (anonymous):

Thank you! :)